A random variable X has the following probability distribution: x 0 1 3 4 5 6 7 P X x ( )  0 k 2k 3k k^2 2k^2 7k^2+ k  A. Find k. 2 B. Evaluate P (1<=X< 6).


1
Expert's answer
2021-06-07T01:42:29-0400

We can find "k" from equality: "1=\sum_{k=0}^7P(X=k)=k+2k+3k+k^2+2k^2+7k^2+k=7k+10k^2" . We solve: "10k^2+7k-1=0". "k=\frac{-7\pm\sqrt{89}}{20}." "k_1\approx-0.822," "k_2\approx0.122".


"P(X\leq1<6)=P(X=1)+P(X=2)+P(X=3)+P(X=4)+P(X=5)=k+2k+3k+k^2+2k^2=6k+3k^2"